The president of Luna Sea Corporation is concened that the year-end inventory of the company is less than it should be. A physical count was made at year-end and the costing was accurately calculated using FIFO. The president asks you to estimae the year-end inventory from the following information: 1. Retail inventory method At retail At c should be Sales Less: Sales returns Net sales Cost of goods sold Opening inventory Purchases Less: Purchases returns Transportation-in Cost of goods available for sale Ending inventory Cost of goods sold Gross proft At retoil Atcost $200,000 Sales Sales returns and allowances 5,000 214,500 $102,224 526 4,000 21,053 Purchases returns and 1,000 Opening inventory 40,000 Required: 1 Calculate the estimated ending inventory at cost using the retail inventory method Mark-up rate 190% 2 Estimated inventory lost Estimated ending inventory Actual inventory on hand (given) 2 Calculate the amount of inventory discrepancy at cost assuming: Actual ending inventory, at cost, is $28,634 Estimated inventory lost
Answer #1
1 At retail At cost,should be
Sales 200000 100220
(200000*50.11%)
Less: Sales returns 5000 2505.5
(5000*50.11%)
Net sales 195000 97714.5
(195000*50.11%)
Cost of goods sold:
Opening inventory 40000 21053
Purchases 214500 102224
Less: Purchase returns -1000 -256
Add: Transportation-in 4000
Cost of goods available for sale 253500 127021
Ending inventory 29306.5
Cost of goods sold 97714.5
(Net sales at cost)
Gross profit 97285.5
Note:1
Cost to retail percetage=Cost of goods avilable for sale at cost/Cost of goods available for sale at retail=127021/253500=0.5011=50.11%
Ending inventory=cost of goods available for sale at cost-Net sales at cost=1270121-97714.50=$ 29306.50
Gross profit=Net sales at retail-cost of goods sold=195000-97714.5=$ 97285.5
2 Estimated inventory lost:
$
Estimated ending inventory 29306.5
Less: Actual inventory on hand (given) 28634
Estimated inventory lost 672.5